### v1

# Dialect Matters: Cross-Lingual ASR Transfer for Low-Resource Indic Language Varieties

Abstract
We conduct an empirical study of cross-lingual transfer using spontaneous, noisy, and code-mixed speech across a wide range of Indic dialects and language varieties. Our results indicate that although ASR performance is generally improved with reduced phylogenetic distance between languages, this factor alone does not fully explain performance in dialectal settings. Often, fine-tuning on smaller amounts of dialectal data yields performance comparable to fine-tuning on larger amounts of phylogenetically-related, high-resource standardized languages. We also present a case study on Garhwali, a low-resource Pahari language variety, and evaluate multiple contemporary ASR models. Finally, we analyze transcription errors to examine bias toward pre-training languages, providing additional insight into challenges faced by ASR systems on dialectal and non-standardized speech.
[ Script=Devanagari, Scale=MatchLowercase, Path=./, ]
Dialect Matters: Cross-Lingual ASR Transfer
for Low-Resource Indic Language Varieties
Akriti Dhasmana  and Aarohi Srivastava  and David Chiang Computer Science and Engineering University of Notre Dame Notre Dame, IN, USA {adhasman, asrivas2, dchiang} @nd.edu
1 Introduction
Automatic speech recognition (ASR) systems for Indic languages have made significant progress through large-scale multilingual pre-training on high-resource, standardized languages. However, these advances have largely overlooked the linguistic reality of the region, where speech is characterized by extensive dialectal variation, spontaneous and noisy recording conditions, and frequent code-switching. As a result, state-of-the-art Indic ASR models often perform poorly when applied to low-resource dialects and language varieties, even those that are closely related to languages seen during pre-training.
Such efforts for Indic ASR typically rely on pre-training on a small set of mainstream languages such as Hindi, Marathi, and Bengali, under the assumption that phylogenetic similarity will facilitate transfer to related varieties. While this assumption provides a useful baseline in many cross-lingual settings, its sufficiency for dialectal speech remains underexplored. In practice, dialects and non-standard varieties often exhibit distinct phonological, lexical, and orthographic properties that are not well-captured by standard language data, raising questions about the effectiveness of relying solely on high-resource languages for transfer in dialectal ASR.
In this work, we examine cross-lingual transfer for Devanagari-script Indic varieties using spontaneous, noisy, and code-mixed speech. Rather than focusing solely on phylogenetic similarity, we investigate how different choices of fine-tuning data affect ASR performance on dialectal speech. Our results suggest that incorporating even limited amounts of dialectal speech during fine-tuning can be as effective as, or more effective than, relying on larger amounts of standardized language data, highlighting the role of dialectal variation in shaping transfer behavior.
To complement our cross-lingual analysis, we present a case study on Garhwali, a low-resource Pahari language variety that has been largely absent from prior ASR research. We evaluate several self-supervised speech models on Garhwali and perform a detailed error analysis to characterize the challenges posed by dialectal variation, code-mixing, and model bias toward pre-training languages. This case study provides concrete insights into the limitations of current Indic ASR models when applied to low-resource language varieties.
Contributions
In this paper, we make three contributions:
  1. 1.
    We provide a comprehensive empirical analysis of cross-lingual transfer for low-resource Devanagari-script Indic dialects and show that fine-tuning on other dialects is often more effective than relying on closely-related high-resource standardized languages.
  2. 2.
    We present the first detailed ASR study for Garhwali, including model evaluation and qualitative and quantitative error analysis.
  3. 3.
    We introduce a general framework for quantifying bias toward pre-training languages in dialectal ASR, enabling systematic analysis of model behavior on non-standardized and code-mixed speech.
2 Related Work
Recent work has investigated improving speech recognition for Indic dialects and language varieties. For instance, Kumar et al. (2025) evaluate a multilingual dialect identification and ASR system that combines Conformer-based speech representations with a RoBERTa model across 33 dialects of 8 Indic languages using read speech. Related efforts have also aimed to develop comprehensive ASR toolkits for Indic languages Singh Chadha et al. (2022). A similar study on standard to dialectal transfer learning in German language has been done Blaschke et al. (2025). However, cross-dialectal ASR adaptation across Indic languages, particularly in zero-shot settings and on spontaneous speech, has not yet been systematically studied, which we address in this work.
The effects of orthographic irregularity and transcription variability on ASR performance have been examined in other language contexts, including cross-family settings Taguchi and Chiang (2024) and Swiss German dialectal ASR Nigmatulina et al. (2020). We build on these findings to analyze how similar forms of orthographic variation impact ASR performance for Indic dialects.
Prior work on speech technology for Garhwali has primarily focused on language identification Gusain et al. (2023) and the creation of domain-specific datasets, such as for agriculture-related applications Riyal et al. (2016). To the best of our knowledge, no prior work has involved training or evaluating an ASR model for Garhwali.
The widespread presence of multiple language varieties in India, shaped by historical and sociolinguistic factors, has led to extensive code-mixing in everyday speech, posing additional challenges for ASR systems. While code-mixed speech has been studied in specific settings such as Hindi-Marathi ASR Palivela et al. (2025), we propose a framework for quantifying transcription errors arising from code-mixing in dialectal ASR.
3 Languages and Data
The Indic linguistic landscape is characterized by a high degree of diversity, encompassing hundreds of languages and dialects with varying levels of standardization and resource availability. Many widely spoken languages coexist with numerous regional dialects that differ substantially in phonology, morphology, and lexicon, despite phylogenetic relatedness. In everyday use, speakers frequently engage in code-mixing, particularly with English, and in spontaneous and acoustically noisy environments. These factors pose major challenges for ASR systems trained primarily on clean, standardized, and monolingual speech. We present a phylogenetic tree of the languages included in our experiments in Figure 1.
Most existing ASR resources for Indic languages focus on a small subset of high-resource, standardized languages, leaving dialectal varieties underrepresented or entirely absent. To address this gap, we fine-tune and evaluate our models on the VAANI dataset (VAANI, 2025). [^1] https://huggingface.co/datasets/ARTPARK-IISc/VAANI VAANI captures the rich linguistic diversity present across speakers of Indic languages, including different accents, grammatical variation, loan words, and code-switching patterns, all of which are vital to include in the context of Indic ASR.
VAANI consists of spontaneous speech collected by prompting participants to describe images in their local dialect. The dataset contains over 150,000 hours of speech, of which approximately 10% is transcribed, and covers speakers from 773 districts across India. As Indic languages are written in multiple scripts, we focus our analysis on language varieties written in the Devanagari script. Due to the broad coverage of dialects and language varieties, the amount of available data varies substantially across languages in the dataset; we account for this variability as much as possible in our experimental design.
Indo-Aryan LanguagesContinental Indo-Aryan LanguagesIndo-Aryan Eastern ZoneHalbiWER 0.87Midlands $Indo-Aryan$ Northern zoneCentral PahariKumaoniWER 0.60GarhwaliWER 0.82Eastern PahariNepaliWER 0.96ShaurasenicBihari GroupBhojpuriWER 0.73MaithiliWER 0.70BajjikaWER 0.78AngikaWER 0.73MagadhiWER 0.70SurjapuriWER 0.88ThethiWER 0.71SadriWER 0.74Eastern HindiChhattisgarhiWER 0.76SurgujiaWER 0.75AwadhiWER 0.77Central ZoneWestern HindiHindiWER 0.50KhariboliWER 0.61HaryanviWER 0.77BundeliWER 0.65ApabhramsicGujarati-Rajasthani GroupRajasthani GroupRajasthaniWER 0.76MarwariWER 0.74JaipuriWER 0.53Indo-Aryan Southern ZoneMarathicMarathi-KonkaniOld-Modern-MarathiMarathiWER 0.87Goan KonkaniKonkaniWER 0.92MalvaniWER 0.88
Figure 1: Subset of the Indo-European language family tree showing Devanagari-script Indic languages in the VAANI dataset, based on Glottolog Hammarström et al. (2024). Languages are annotated with WER for IndicWav2Vec-Hindi. Blue highlights indicate languages used during pre-training.
Beyond conducting cross-lingual experiments across Indic language varieties, we also conduct a detailed analysis of ASR for one nonstandard language variety. We select Garhwali for this case study. [^2] According to the Post-1971 Census of India, these languages were recognized as “mother tongues” and were designated as dialects of Hindi Khubchandani (1991); however, some linguists argue that Garhwali and other Pahari languages are distinct languages Gusain et al. (2023). Garhwali is in the Pahari subgroup [^3] We refer to the lowest non-leaf nodes in a phylogenetic tree as subgroups. spoken in the Himalayan region, particularly in the state of Uttarakhand Gusain et al. (2023).
4 Experiments and Results
In our experiments, we aim to answer the following questions:
  1. RQ1.
    How does a strong Indic ASR baseline perform across a diverse set of low-resource language varieties, particularly those not seen during model pre-training?
  2. RQ2.
    To what extent does orthographic variability correlate with ASR performance in Indic language varieties?
  3. RQ3.
    How does cross-lingual transfer for Indic ASR vary with phylogenetic distance, and does this relationship hold for dialectal speech?
In addition, we conduct a focused case study on Garhwali to characterize systematic error patterns in dialectal ASR. We analyze how systemic bias towards pre-training languages (Hindi) manifest in transcriptions generated by ASR models fine-tuned on a dialect (Garhwali).
Metrics
We employ two standard metrics, word error rate (WER) and character error rate (CER), to evaluate the generated transcriptions. Both metrics are calculated based on Levenshtein distance, representing the minimum number of insertions, deletions, and substitutions required to align the hypothesis with the reference text. We report WER in our main results and include CER in the appendix.
4.1 Analyzing Broader Patterns in Indic ASR
RQ1: Baseline Assessment
We begin by assessing the performance of a state-of-the-art Indic ASR model, IndicWav2Vec Kumar et al. (2022), on language varieties present in the VAANI dataset. IndicWav2Vec is based on the Wav2Vec 2.0 architecture Baevski et al. (2020), and is pre-trained on 17,000 hours of unlabeled clean speech data from YouTube, as well as Newsonair data curated from radio channels, across 40 Indic languages. In our experiments, we employ IndicWav2Vec-Hindi [^4] https://huggingface.co/ai4bharat/IndicWav2Vec-Hindindi (IndicWav2Vec fine-tuned on Hindi), since pre-trained ASR models largely learn language-agnostic acoustic representations during pre-training and require language-specific fine-tuning to map these representations to text transcriptions.
Figure 2: Hapax legomena (once-seen word types) % per language in the test split of VAANI plotted against WER (Pearson’s $\rho=0.705$, $p=4\times 10^{-4}$) and CER (not significant) using IndicWav2Vec-Hindi.
Results
We evaluate IndicWav2Vec-Hindi on test samples for all 30 Devanagari-script varieties in VAANI (up to 1 hour each) and report these results as part of a phylogenetic tree in Figure 1. We also divide the results into two tables depending on whether the language was used to pre-train IndicWav2Vec (Table 7) or it was not explicitly seen by the model (Table 8). These results show that, despite being fine-tuned on Hindi, IndicWav2Vec-Hindi achieves a best-case word error rate of 50.4% on VAANI Hindi test speech (one hour). Similarly high error rates are observed for several languages seen during pre-training, indicating that pre-training alone does not ensure robust performance on spontaneous and noisy speech, even for languages included in the pre-training corpus. We also observe substantial variation in performance across dialects and closely-related language varieties. These observations suggest that other factors, such as geographical proximity or phonological feature similarity, could have an impact on performance, motivating our subsequent analyses.
RQ2: Orthographic Consistency
Orthographic consistency is an important factor in achieving high ASR performance, as inconsistent spellings introduce additional variability that can increase modeling and decoding errors Nigmatulina et al. (2020). This challenge is particularly pronounced for Indic languages that encompass multiple dialects where orthographic conventions are not fixed. To quantify orthographic irregularity in the data, we measure the frequency of each word type in the transcripts and analyze the proportion of tokens that occur only once, along with the type-to-token ratio. We then examine how these measures correlate with ASR performance.
Results
We observe a clear trend between orthographic variability, as measured by the proportion of hapax legomena (unique words) in the test set, and ASR performance. Figure 2 shows that languages with a higher percentage of unique words exhibit higher WERs when evaluated with IndicWav2Vec-Hindi (Pearson’s $\rho=0.705$, $p=4\times 10^{-4}$), indicating increased difficulty for languages with less consistent orthographic conventions (e.g., Thethi, Surjapuri). A similar but weaker trend is observed for CER.
Additional evidence is provided by the type-token statistics reported in Table 9. Several languages (e.g., Gondi, Thethi) exhibit a high number of distinct word types relative to the total number of tokens, resulting in elevated type-to-token ratios. These measures reflect greater orthographic and lexical variability and are associated with higher ASR error rates. Taken together, these results suggest that orthographic inconsistency is an important contributing factor to reduced ASR performance in nonstandard Indic varieties.
Figure 3: Cross-Lingual Performance of w2vBERT models fine-tuned on 1 to 7 hours of data per language vs. off-the-shelf IndicWav2Vec-Hindi.
RQ3: Cross-Lingual Transfer
We wanted to examine whether there are any performance gains from cross-dialectal fine-tuning in zero-shot settings. For each of these experiments, we fine-tune w2v-bert-2.0555https://huggingface.co/facebook/w2v-bert-2.0 (selected through preliminary comparison of models) on 1 to 7 hours of speech, depending on the availability. Each model was then evaluated on the test set of all Devanagari script languages available from VAANI. We compute correlations between the error rate and phylogenetic distance under two conditions: across all evaluation languages and restricted to nonstandard test varieties. We additionally measure the association between the error rate and whether the fine-tuning language represents dialectal or standard speech. We consider the following languages from the dataset to be dialects or nonstandard varieties: Angika Awadhi, Bajjika, Bhili, Chhattisgarhi, Garhwali, Halbi, Haryanvi, Jaipuri, Khariboli, Khortha, Kumaoni, Magadhi, Malvani, Marwari, Nagpuri, Surjapuri, and Thethi. [^6] None of these languages belong to the Eighth Schedule to the Indian Constitution listing the officially recognized languages (https://en.wikipedia.org/wiki/Eighth_Schedule_to_the_Constitution_of_India).
Results
Figure 3 summarizes cross-lingual transfer performance for w2vBERT models fine-tuned on 1 to 7 hours of data per language and evaluated in a zero-shot setting across a wide range of Indic language varieties, providing a holistic view of how ASR performance varies as a function of both the fine-tuning and the evaluation languages.
Cross-Lingual Transfer Trends
Across all evaluation languages, we observe a clear association between phylogenetic distance and ASR performance: larger distances between the fine-tuning and evaluation languages generally correspond to higher word error rates (Spearman’s $\rho=0.333$, $p=1.11\times 10^{-7}$). This result aligns with prior findings in cross-lingual ASR and confirms that phylogenetic relatedness provides a useful baseline for transfer across Indic languages. However, this aggregate trend does not fully characterize model behavior when evaluation is restricted to dialectal speech.
Dialectal Transfer Effects
When focusing specifically on dialects and nonstandard test varieties, the heatmap reveals additional structure that is not explained by phylogenetic distance alone. Although the association between phylogenetic distance and word error rate remains statistically significant in this setting (Spearman’s $\rho=0.274$, $p=3.604\times 10^{-4}$), several consistent deviations from this trend are observed. In particular, models fine-tuned on nonstandard varieties frequently outperform models fine-tuned on phylogenetically closer but standard, higher-resource languages.
One representative example from Figure 3 is the strong transfer from Marwari (a dialect of Rajasthani) to Kumaoni (a Pahari language variety), even though the two belong to distinct phylogenetic subgroups and are geographical distant. Another example is the consistently competitive performance of the models fine-tuned on Marwari and Magadhi across multiple evaluation language varieties, often outperforming models fine-tuned on higher-resource standardized languages such as Hindi, Marathi, and Rajasthani. Notably, this trend holds despite the fact that the Magadhi model is trained on less data (5 vs. 7 hours). These cases illustrate that while fine-tuning on a closely related high-resource standardized language may be the most natural strategy, it is not necessarily the most effective choice for transfer to dialects and related language varieties.
To further examine this pattern, we analyze whether the dialectal status of the fine-tuning language itself is associated with performance on unseen dialects. We find a statistically significant trend indicating that fine-tuning on dialectal speech is associated with lower WER on dialectal evaluation sets (point-biserial correlation $r_{pb}=-0.196$, $p=5.79\times 10^{-3}$). This trend holds even when the available dialectal training data is smaller than that of the corresponding standardized languages, suggesting that dialectal fine-tuning captures information that is not adequately represented by mainstream language data alone.
Controlled In-Group Transfer Experiments
Fine-Tuning Lang Hours Jaipuri
WER CER
Rajasthani + Marwari 16.8 0.309 0.151
Marwari only (dialect) 17.2 0.301 0.162
Table 1: Effect of fine-tuning language choice on Jaipuri ASR (Rajasthani subgroup). Mainstream language is in bold.
Fine-Tuning Lang Hours Thethi
WER CER
Angika, Bajjika, Bhojpuri, Khortha, Magadhi, Maithili, Sadri 50.6 0.412 0.132
Angika, Bajjika, Khortha, Magadhi, Sadri (dialects only) 15.6 0.353 0.147
Table 2: Effect of mainstream language inclusion on Thethi ASR (Bihari subgroup). Mainstream language is in bold.
To isolate the effect of dialectal vs. standard fine-tuning data from broader cross-lingual trends, we conduct controlled transfer experiments within individual phylogenetic subgroups. These experiments compare models fine-tuned on dialectal varieties against models fine-tuned on higher-resource standardized languages within the same subgroup.
Fine-Tuning Lang Hours Haryvani
WER CER
Hindi + Bundeli + Khariboli 101.7 0.930 1.452
Bundeli + Khariboli (dialects only) 101.7 0.513 0.267
Table 3: Impact of fine-tuning (FT) languages on Haryanvi ASR (Western Hindi subgroup). Mainstream language is in bold.
Across the Rajasthani, Western Hindi, and Bihari subgroups, we consistently observe that fine-tuning on small amounts of dialectal speech yields performance comparable to or better than fine-tuning on substantially larger amounts of standardized language data. For example, within the Rajasthani subgroup (Table 1), a model fine-tuned solely on a dialectal variety (Marwari) achieves comparable performance on an unseen dialect (Jaipuri) to a model trained on a combination of dialectal and macro-language data. Similar trends are observed in the Western Hindi and Bihari subgroups (Tables 3 and 2), where excluding higher-resource mainstream languages does not degrade, and in some cases improves, performance on dialectal evaluation sets.
Taken together, the heatmap analysis and controlled in-group experiments demonstrate that while phylogenetic similarity provides a useful starting point for cross-lingual transfer, effective ASR for dialectal speech depends critically on the inclusion of dialectal training data itself.
4.2 Garhwali ASR Case Study
Baseline Selection
In order to determine the ideal model architecture for Garhwali, we compared several widely-used ASR models:
  • Wav2Vec2 Baevski et al. (2020)
  • HuBERT Hsu et al. (2021)
  • XLS-R Babu et al. (2022)
  • Whisper Radford et al. (2023)
  • w2vBERT Chung et al. (2021)
We fine-tune each of these models on the Garhwali subset of VAANI, using an 87%–6%–7% train–test–validation ratio.
We fine-tuned and evaluated several self-supervised speech models on the Garhwali subset of the VAANI dataset. Table 4 reports performance across these models.
Results
Among the evaluated configurations, the w2vBERT-based model achieves the lowest error rates and is therefore selected for subsequent analysis. We compare this model against IndicWav2Vec-Hindi (used in our cross-lingual experiments). Our fine-tuned model performs better than the IndicWav2Vec-Hindi model; however, the resulting error rates remain high, indicating that Garhwali ASR remains challenging even with dialect-specific fine-tuning. We therefore focus our error analysis on this best-performing configuration to better understand the remaining sources of error.
Model Metric Optimized WER CER
XLS-R WER 0.650 0.270
wav2vec2-BERT CER 0.493 0.193
Whisper-small CER 0.629 0.650
HuBERT CER 0.515 0.199
Table 4: Comparison of speech models fine-tuned on Garhwali. Training details are in Table 10.
Error Analysis
We conduct an extensive error analysis of the best fine-tuned model on Garhwali to identify the following:
Inconsistent English Transliteration
[…]
This variability introduces ambiguity during training and evaluation, complicating the model’s ability to learn consistent lexical representations. Whether standardizing transliterated forms would improve dialectal ASR performance or instead remove linguistically meaningful variation remains an open question.
Model # Non-Hi Correct To Hi To Wrong
wav2vec2-BERT 1873 34.8% 23.3% 38.8%
Indicw2v2 2109 4.2% 37.3% 42.9%
Table 5: Comparison of w2vBERT model fine-tuned on Garhwali against IndicWav2Vec-Hindi on Garhwali non-Hindi word handling. Key: Hi = Hindi.
Bias Towards Hindi from Pre-training
We identify non-Hindi words in the ground-truth transcripts using Hindi-HunSpell, [^7] https://github.com/Shreeshrii/hindi-hunspell a Hindi spell-checker. These includes Garhwali words, as well as transliterated English words; we separate the English words (see below). We then assess which non-Hindi words are preserved in the generated transcription vs. which are transcribed to a Hindi word. We analyze commonly omitted, added, and substituted characters in the generated transcriptions, qualitatively identifying systematic errors and highlighting areas for future improvement. We also observe systematic bias towards Hindi arising from model pre-training. Many errors involve Garhwali words or transliterated English terms being incorrectly normalized to valid Hindi words.
[…]
As shown in Table 5, the w2vBERT model fine-tuned on Garhwali converts approximately 22.9% of non-Hindi terms into Hindi words, while preserving roughly one-third of such terms. In contrast, IndicWav2Vec-Hindi preserves fewer than 5% of non-Hindi words, converting or distorting the majority. These results indicate that ASR pre-training on mainstream Indic languages can substantially hamper a model’s ability to retain dialect-specific information.
Actual Predicted # Actual Predicted #
Deva IPA Deva IPA Deva IPA Deva IPA
[…]
Table 6: Most frequent transcription errors for Garhwali ASR. Key: ␣ = space; $\epsilon$ = empty string; Deva = Devanagari script.
Character-Level Error Patterns
[…]
5 Conclusion
In this work, we present a comprehensive study of dialectal speech recognition across a diverse set of Devanagari-script Indic language varieties, with a particular focus on understanding how dialectal variation interacts with cross-lingual transfer in low-resource ASR. Through a large-scale empirical evaluation, we find that while ASR performance is generally associated with phylogenetic distance across languages, this factor alone does not explain performance in the dialectal setting. In particular, when evaluating on dialects in the zero-shot setting, we observe lower word error rates when the fine-tuning language is a dialect or nonstandard variety. In many cases, fine-tuning on small amounts of dialectal speech yields performance comparable to or better than fine-tuning on larger amounts of phylogenetically closer, high-resource standardized languages.
Across multiple phylogenetic subgroups, our results consistently demonstrate that including higher-resource mainstream languages during fine-tuning does not reliably improve zero-shot ASR performance on dialectal evaluation sets. Instead, whether the fine-tuning data itself reflects dialectal speech emerges as a more informative predictor of performance than phylogenetic proximity alone. These findings highlight the importance of treating dialects as distinct acoustic and linguistic entities rather than as minor variants of standardized languages when designing ASR systems.
We further present the first detailed ASR analysis for Garhwali, a nonstandard Pahari language variety, and show that a w2vBERT-based model fine-tuned on Garhwali achieves the best performance among the evaluated architectures. Although the resulting word error rate of 49.3% remains insufficient for fully automated transcription, this case study illustrates both the challenges of dialectal ASR and the benefits of dialect-specific modeling.
Finally, we introduce a quantitative framework for analyzing bias toward pre-training languages in dialectal ASR. Our error analysis reveals substantial bias toward Hindi in both multilingual and Hindi-fine-tuned models, manifesting in systematic normalization of dialectal and code-mixed forms. This framework underscores the need for dialect-aware data selection and modeling strategies in future ASR systems.
Overall, our findings suggest that effective ASR for low-resource dialects requires moving beyond default assumptions of phylogenetic similarity and toward evaluation and modeling practices that explicitly account for dialectal variation.
Limitations
Our study is based on the VAANI dataset, which contains varying amounts of data across language varieties. Although this diversity allows us to evaluate ASR performance across a wide range of realistic dialectal settings, differences in dataset size may influence performance comparisons across languages. We mitigate this effect where possible by controlling the amount of fine-tuning data used across languages, though some variability remains inherent to the dataset.
In addition, VAANI consists of spontaneous and naturally-occurring speech collected across diverse regions. While this enables evaluation under realistic acoustic and conversational conditions, the presence of background noise, disfluencies, and region-specific recording environments may introduce additional variability in model performance.
Our analysis is limited to Indic dialects and language varieties written in the Devanagari script. Although this choice allows for controlled comparisons within a shared orthographic system, it excludes Indic languages written in other scripts, and our findings may not directly generalize beyond the Devanagari-script subset.
Finally, our experiments focus on a specific set of self-supervised ASR architectures and fine-tuning strategies. While these models are representative of widely used contemporary approaches, different architectures or training objectives may exhibit different transfer behaviors.
Appendix A Appendix
A.1 Baseline Evaluation
We used the IndicWav2Vec-Hindi model to establish a baseline for the performance of off-the-shelf models on the languages in the VAANI dataset. We further divided the results into two tables based on whether the Evaluation Language was seen by the IndicWav2Vec-Hindi model during pretraining or not.
Language CER WER
Bhojpuri 0.786 0.732
Chhattisgarhi 0.783 0.757
Hindi 0.752 0.504
Konkani 0.851 0.916
Maithili 0.763 0.696
Marathi 0.821 0.870
Nepali 0.785 0.957
Rajasthani 0.768 0.762
Table 7: IndicWav2Vec-Hindi performance on languages seen during pre-training.
Language CER WER Language CER WER
Angika 0.791 0.725 Khortha 0.829 0.743
Awadhi 0.774 0.771 Kumaoni 0.759 0.596
Bajjika 0.784 0.784 Kurukh 0.825 0.875
Bhili 0.860 0.944 Magadhi 0.782 0.700
Bundeli 0.774 0.646 Malvani 0.808 0.881
Garhwali 0.760 0.820 Marwari 0.752 0.740
Gondi 0.718 0.945 Nagpuri 0.797 0.834
Halbi 0.788 0.873 Sadri 0.742 0.738
Haryanvi 0.811 0.766 Surgujia 0.762 0.747
Jaipuri 0.696 0.532 Surjapuri 0.829 0.880
Khariboli 0.784 0.605 Thethi 0.796 0.707
Table 8: IndicWav2Vec-Hindi performance on related but unseen languages.
A.2 Orthographic Analysis
We computed the total number of words (tokens), total number of unique words (types) and the total number of words that only appear once (Hapax) per all the languages in the VAANI dataset. We also compute the total percentage of words that occur only once (Hapax%) and the ratio of unique words to total number of words (TTR) in the table 9.
Language Tokens Types Hapax Hapax % TTR
Gondi 3,136 1,562 1,168 74.78 0.4981
Thethi 1,556 654 469 71.71 0.4203
Bhili 1,420 632 448 70.89 0.4451
Kurukh 2,022 798 559 70.05 0.3947
Nepali 2,169 776 536 69.07 0.3578
Surjapuri 2,450 860 593 68.95 0.3510
Sadri 6,599 2,096 1,417 67.60 0.3176
Nagpuri 1,571 612 404 66.01 0.3896
Malvani 9,123 2,501 1,636 65.41 0.2741
Konkani 31,668 6,327 4,055 64.09 0.1998
Awadhi 2,346 887 561 63.25 0.3781
Surgujia 5,611 1,524 950 62.34 0.2716
Bundeli 7,734 1,613 969 60.07 0.2086
Khariboli 15,345 2,465 1,476 59.88 0.1606
Angika 34,348 5,420 3,233 59.65 0.1578
Garhwali 77,030 10,431 6,216 59.59 0.1354
Haryanvi 3,747 983 584 59.41 0.2623
Halbi 16,716 3,351 1,979 59.06 0.2005
Marathi 171,424 16,876 9,919 58.78 0.0984
Bajjika 32,049 4,681 2,717 58.04 0.1461
Magadhi 55,891 7,101 4,093 57.64 0.1271
Khortha 38,324 4,499 2,584 57.43 0.1174
Jaipuri 2,074 599 339 56.59 0.2888
Chhattisgarhi 169,861 13,024 7,287 55.95 0.0767
Bhojpuri 209,523 15,243 8,508 55.82 0.0728
Marwari 85,883 8,279 4,582 55.34 0.0964
Kumaoni 22,911 3,190 1,737 54.45 0.1392
Maithili 194,541 14,317 7,779 54.33 0.0736
Hindi 156,389 8,156 4,381 53.72 0.0522
Rajasthani 116,180 7,988 4,273 53.49 0.0688
Table 9: Hapax-Legomena (unique words) per training set.
A.3 Garhwali ASR
We evaluated multiple model architecture varieties on the Garhwali language test split from the VAANI dataset. For each model architecture, we first evaluated the performance without finetuning first and then after finetuning on the Training split of the Garhwali dataset. For each finetuning experiment except for the Whisper model, we generated the vocab of characters present in the training and validation split.
Model Type Tr. Lang FT Lang Gen Vocab? Metric Tr. Error Val Loss Tr. Loss Steps Test WER Test CER
wav2vec2CTC - Garhwali Yes WER 0.769 1.576 0.403 6500 0.769 -
wav2vec2CTC - - Yes - - - - - 1.015 2.436
XLS-R - Garhwali Yes WER 0.735 1.554 0.171 1600 0.650 0.270
XLS-R - - Yes - - - - - 1.000 1.292
wav2vec2 BERT - Garhwali Yes CER 0.197 0.940 0.397 1200 0.493 0.193
wav2vec2 BERT - - Yes - - - - - 1.000 1.600
Whisper-small Hindi Garhwali No CER 0.242 0.974 0.001 4000 0.629 0.650
Whisper-small Hindi - No - - - - - 2.518 1.262
HuBERT - Garhwali Yes CER - - - - 0.515 0.199
Table 10: Comparison of Speech Models (Garhwali Fine-tuning vs. Baselines)

### v2

@@
   3. 3.
-    We introduce a general framework for quantifying bias toward pre-training languages in dialectal ASR, enabling systematic analysis of model behavior on non-standardized and code-mixed speech.
+    We introduce a diagnostic approach for quantifying bias toward pre-training languages in dialectal ASR, enabling systematic analysis of model behavior on non-standardized and code-mixed speech.
 2 Related Work
-Recent work has investigated improving speech recognition for Indic dialects and language varieties.
-For instance, Kumar et al.
-(2025) evaluate a multilingual dialect identification and ASR system that combines Conformer-based speech representations with a RoBERTa model across 33 dialects of 8 Indic languages using read speech.
-Related efforts have also aimed to develop comprehensive ASR toolkits for Indic languages Singh Chadha et al.
+Recent work has highlighted systematic performance gaps between standard language varieties and regional or minority dialects across NLP and speech technologies.
+Kantharuban et al.
+(2023) provide a large-scale evaluation of state-of-the-art models for machine translation and automatic speech recognition across regional dialects of multiple high- and low-resource languages, showing that dialectal performance disparities are widespread and variably correlated with linguistic, social, and data-related factors.
+Complementary to this line of work, Blaschke et al.
+(2025) study standard-to-dialect transfer in spoken and written settings for German, demonstrating that speech-based models are more robust to dialectal variation than text-based or cascaded approaches, particularly when orthographic normalization is involved.
+In the Indic context, prior work has investigated dialectal variation primarily through multilingual or language-specific ASR systems.
+For example, Kumar et al.
+(2025) evaluate a multilingual dialect identification and ASR pipeline across 33 dialects of eight Indic languages using read speech, while other efforts have focused on building comprehensive ASR toolkits for standardized Indic languages Singh Chadha et al.
 (2022).
-A similar study on standard to dialectal transfer learning in German language has been done Blaschke et al.
-(2025).
-However, cross-dialectal ASR adaptation across Indic languages, particularly in zero-shot settings and on spontaneous speech, has not yet been systematically studied, which we address in this work.
+However, cross-dialectal and cross-lingual transfer for Indic ASR—especially in zero-shot settings and on spontaneous, non-standardized speech—has not yet been systematically examined.
+Our work addresses this gap by analyzing cross-lingual transfer behavior and dialectal bias in low-resource Indic language varieties.
 The effects of orthographic irregularity and transcription variability on ASR performance have been examined in other language contexts, including cross-family settings Taguchi and Chiang (2024) and Swiss German dialectal ASR Nigmatulina et al.
@@
 While code-mixed speech has been studied in specific settings such as Hindi-Marathi ASR Palivela et al.
-(2025), we propose a framework for quantifying transcription errors arising from code-mixing in dialectal ASR.
+(2025), we provide a systematic way to quantify transcription errors arising from code-mixing in dialectal ASR.
 3 Languages and Data
 The Indic linguistic landscape is characterized by a high degree of diversity, encompassing hundreds of languages and dialects with varying levels of standardization and resource availability.
-Many widely spoken languages coexist with numerous regional dialects that differ substantially in phonology, morphology, and lexicon, despite phylogenetic relatedness.
+Many widely spoken languages coexist with numerous regional dialects that differ substantially in phonology, morphology, and lexicon, despite phylogenetic relatedness Massica (1993).
 In everyday use, speakers frequently engage in code-mixing, particularly with English, and in spontaneous and acoustically noisy environments.
-These factors pose major challenges for ASR systems trained primarily on clean, standardized, and monolingual speech.
+These factors pose major challenges for ASR systems trained primarily on clean, standardized, and monolingual speech Diwan et al.
+(2021).
 We present a phylogenetic tree of the languages included in our experiments in Figure 1.
@@
 VAANI consists of spontaneous speech collected by prompting participants to describe images in their local dialect.
-The dataset contains over 150,000 hours of speech, of which approximately 10% is transcribed, and covers speakers from 773 districts across India.
+The dataset contains over 150,000 hours of speech, of which approximately 10% is transcribed, and covers 156,534 speakers from 773 districts across India.
 As Indic languages are written in multiple scripts, we focus our analysis on language varieties written in the Devanagari script.
@@
 (2023).
-Garhwali is in the Pahari subgroup [^3] We refer to the lowest non-leaf nodes in a phylogenetic tree as subgroups.
+Garhwali is a language that belongs to the Pahari subgroup  [^3] We refer to the lowest non-leaf nodes in a phylogenetic tree as subgroups.
 spoken in the Himalayan region, particularly in the state of Uttarakhand Gusain et al.
@@
 4 Experiments and Results
-In our experiments, we aim to answer the following questions:
+In our experiments, we aim to answer the following questions to analyze broader patterns in Indic ASR:
   1. RQ1.
@@
 We report WER in our main results and include CER in the appendix.
-4.1 Analyzing Broader Patterns in Indic ASR
-RQ1: Baseline Assessment
+4.1 RQ1: Baseline Assessment for Indic ASR
+Setup
 We begin by assessing the performance of a state-of-the-art Indic ASR model, IndicWav2Vec Kumar et al.
@@
 These observations suggest that other factors, such as geographical proximity or phonological feature similarity, could have an impact on performance, motivating our subsequent analyses.
-RQ2: Orthographic Consistency
+4.2 RQ2: Orthographic Consistency
+Setup
 Orthographic consistency is an important factor in achieving high ASR performance, as inconsistent spellings introduce additional variability that can increase modeling and decoding errors Nigmatulina et al.
@@
 Figure 3: Cross-Lingual Performance of w2vBERT models fine-tuned on 1 to 7 hours of data per language vs. off-the-shelf IndicWav2Vec-Hindi.
-RQ3: Cross-Lingual Transfer
+4.3 RQ3: Cross-Lingual Transfer
 We wanted to examine whether there are any performance gains from cross-dialectal fine-tuning in zero-shot settings.
@@
 We additionally measure the association between the error rate and whether the fine-tuning language represents dialectal or standard speech.
-We consider the following languages from the dataset to be dialects or nonstandard varieties: Angika Awadhi, Bajjika, Bhili, Chhattisgarhi, Garhwali, Halbi, Haryanvi, Jaipuri, Khariboli, Khortha, Kumaoni, Magadhi, Malvani, Marwari, Nagpuri, Surjapuri, and Thethi.
+We consider the following languages from the dataset to be dialects or nonstandard varieties: Angika, Awadhi, Bajjika, Bhili, Chhattisgarhi, Garhwali, Halbi, Haryanvi, Jaipuri, Khariboli, Khortha, Kumaoni, Magadhi, Malvani, Marwari, Nagpuri, Surjapuri, and Thethi.
 [^6] None of these languages belong to the Eighth Schedule to the Indian Constitution listing the officially recognized languages (https://en.wikipedia.org/wiki/Eighth_Schedule_to_the_Constitution_of_India).
@@
 Figure 3 summarizes cross-lingual transfer performance for w2vBERT models fine-tuned on 1 to 7 hours of data per language and evaluated in a zero-shot setting across a wide range of Indic language varieties, providing a holistic view of how ASR performance varies as a function of both the fine-tuning and the evaluation languages.
-Cross-Lingual Transfer Trends
 Across all evaluation languages, we observe a clear association between phylogenetic distance and ASR performance: larger distances between the fine-tuning and evaluation languages generally correspond to higher word error rates (Spearman’s $\rho=0.333$, $p=1.11\times 10^{-7}$).
@@
 However, this aggregate trend does not fully characterize model behavior when evaluation is restricted to dialectal speech.
-Dialectal Transfer Effects
+4.3.1 Dialectal Transfer Effects
 When focusing specifically on dialects and nonstandard test varieties, the heatmap reveals additional structure that is not explained by phylogenetic distance alone.
@@
 Controlled In-Group Transfer Experiments
-Fine-Tuning Lang Hours Jaipuri
+Fine-Tuning Language Hours Jaipuri
 WER CER
 Rajasthani + Marwari 16.8 0.309 0.151
-Marwari only (dialect) 17.2 0.301 0.162
+Marwari (dialect only) 17.2 0.301 0.162
 Table 1: Effect of fine-tuning language choice on Jaipuri ASR (Rajasthani subgroup). Mainstream language is in bold.
-Fine-Tuning Lang Hours Thethi
+Fine-Tuning Language Hours Thethi
 WER CER
@@
 These experiments compare models fine-tuned on dialectal varieties against models fine-tuned on higher-resource standardized languages within the same subgroup.
-Fine-Tuning Lang Hours Haryvani
+Fine-Tuning Language Hours Haryvani
 WER CER
@@
 Taken together, the heatmap analysis and controlled in-group experiments demonstrate that while phylogenetic similarity provides a useful starting point for cross-lingual transfer, effective ASR for dialectal speech depends critically on the inclusion of dialectal training data itself.
-4.2 Garhwali ASR Case Study
-Baseline Selection
+5 Garhwali ASR Case Study
+5.1 Baseline Selection
+Setup
 In order to determine the ideal model architecture for Garhwali, we compared several widely-used ASR models:
@@
 Table 4: Comparison of speech models fine-tuned on Garhwali. Training details are in Table 10.
-Error Analysis
+5.2 Error Analysis
 We conduct an extensive error analysis of the best fine-tuned model on Garhwali to identify the following:
@@
 Although the resulting word error rate of 49.3% remains insufficient for fully automated transcription, this case study illustrates both the challenges of dialectal ASR and the benefits of dialect-specific modeling.
-Finally, we introduce a quantitative framework for analyzing bias toward pre-training languages in dialectal ASR.
-Our error analysis reveals substantial bias toward Hindi in both multilingual and Hindi-fine-tuned models, manifesting in systematic normalization of dialectal and code-mixed forms.
-This framework underscores the need for dialect-aware data selection and modeling strategies in future ASR systems.
+Our quantitative error analysis further reveals substantial bias toward Hindi in both multilingual and Hindi-fine-tuned models, manifesting in systematic normalization of dialectal and code-mixed forms, and underscoring the need for dialect-aware data selection and modeling strategies in future ASR systems.
 Overall, our findings suggest that effective ASR for low-resource dialects requires moving beyond default assumptions of phylogenetic similarity and toward evaluation and modeling practices that explicitly account for dialectal variation.